# On orthogonal decompositions of hermitian Higgs bundles

### v1

@@ Abstract.
An hermitian Higgs bundle is a triple $({\mathfrak{E}},h)=(E,\Phi,h)$, where ${\mathfrak{E}}=(E,\Phi)$ is a Higgs bundle [15] and $(E,h)$ is a holomorphic hermitian vector bundle [12]. As it is well known, several results on holomorphic vector bundles can be extended to the Higgs bundles setting. In this article we show that some classical propositions, involving orthogonal decompositions of holomorphic hermitian vector bundles and the second fundamental form of its holomorphic subbundles, can be extended to hermitian Higgs bundles. The extended results concerning orthogonal decompositions have immediate applications in Higgs bundles, and we include them in the form of two corollaries. The extended results concerning the second fundamental form are previously known in literature [2]. Here we revisit these results and include alternative proofs without using local computations.
@@ 1. Introduction
Roughly speaking a Higgs bundle is a holomorphic vector bundle, together with a section of a certain associated holomorphic vector bundle. Since its introduction by Hitchin [9] and Simpson [15, 16], these geometric objects have played an important role in complex geometry as well as in mathematical-physics. For instance, from the viewpoint of geometry, several results on holomorphic vector bundles concerning special metrics and connections can be extended to Higgs bundles [2, 3, 4, 5, 13, 14]. Similarly, from the viewpoint of physics, Higgs bundles are closely related to dimensional reduction. In fact, in the pioneering work [9], Higgs bundles arise as geometric objects associated to Hitchin’s equations (a two dimensional reduction of the self-dual Yang-Mills equations in four dimensions). On the other hand, Hitchin’s equations can also be seen as the lowest dimensional case of another set of equations introduced by Ward [19] and which are a dimensional reduction of Yang-Mills equations in higher dimensions. The equations introduced by Ward are known as $2k$-Hitchin’s equations and can be directly related to Higgs bundles [6]. It is important to mention that from the very beginning, Higgs bundles are associated to another set of equations, usually known as the Hermitian-Yang-Mills equations [15]. Higgs bundles are to HYM equations as holomorphic vector bundles are to Hermitian-Einstein’s equations [12, 17]. Today, Higgs bundles and all the above equations are of major interest in geometry and physics and there exists a very ample literature related to the topic, see for instance [1, 11, 18, 20]. In particular, Higgs bundles are geometric objects of relevance in the foundational work of Kapustin and Witten [10].
@@
This article is organized as follows. In Section 1 we review part of the elementary theory on holomorphic bundles that will be important later on. In particular, we explicitly recall four classical propositions on holomorphic hermitian vector bundles. In this section we also establish part of the notation and terminology that will be used throughout the article. In Section 2 we summarize some basic concepts and properties on Higgs bundles. We also mention two lemmas concerning short exact sequences and invariance conditions on Higgs bundles that will be relevant in the remaining part of the article. In this section we also revisit some propositions on Higgs bundles [2], which are extensions of well known results associated to the second fundamental form of holomorphic subbundles of holomorphic hermitian vector bundles [12]. We include here alternative proofs of these results that do not use local computations. In Section 3 we establish a couple of propositions concerning orthogonal decompositions of hermitian Higgs bundles. In particular, we include proofs of these results that rely heavily on classical results on holomorphic hermitian vector bundles as well as standard definitions on Higgs bundles. Finally, in this section we show some immediate applications on Higgs bundles of the results previously established. More specifically, we show that using the aforementioned lemmas and by introducing the notion of $C^{\infty}$ Higgs orthogonal complement, the extended results on Higgs bundles involving orthogonal decompositions imply two corollaries, which can be established by referring exclusively to objects in the Higgs category.
@@
The main purpose of this article is to establish a couple of propositions as well as to revisit two other results on hermitian Higgs bundles [2]. The first propositions are natural extensions of well known propositions on orthogonal decompositions of holomorphic hermitian vector bundles. These classical results are included in the form of propositions (4.18) and (6.14) in the first chapter of the celebrated textbook of Kobayashi [12]. There are two other classical results, closely related to the previous propositions and which appear in the same reference as propositions (6.4) and (6.6). Extensions to hermitian Higgs bundles of these latter results have been previously studied in [2]. For the reader convenience, in this section we review some elementary definitions as well as summarize all these classical results on holomorphic vector bundles (for more details the reader can see [7, 8, 12]).
@@
Throughout this article, $M$ is an $n$-dimensional compact Kähler manifold and $E$ is a rank $r$ holomorphic vector bundle over it. We denote by $\Omega^{1,0}$ and $\Omega^{0,1}$ the holomorphic cotangent bundle to $M$ and its complex conjugate bundle and by $\Omega^{p,q}$, with $0\leq p,q\leq n$, the $C^{\infty}$ complex vector bundle on $M$ obtained by taking wedge products ($p$ and $q$ times) of $\Omega^{1,0}$ and $\Omega^{0,1}$. If $1\leq s\leq 2n$ we have $\Omega^{s}=\oplus_{p+q=s}\Omega^{p,q}$ and we denote by $A^{s}$ and $A^{p,q}$ (resp. $A^{s}(E)$ and $A^{p,q}(E)$) the spaces of $C^{\infty}$ complex $s$-forms and $(p,q)$-forms on $M$ (resp. forms on $M$ with coefficients in $E$), i.e., they are the spaces of $C^{\infty}$ sections of $\Omega^{s}$ and $\Omega^{p,q}$ (resp. of $\Omega^{s}\otimes E$ and $\Omega^{p,q}\otimes E$). As it is known, any hermitian metric $h$ in $E$ determines a unique connection $D_{h}$ in $E$, usually known as the Chern or the hermitian connection of $h$. Following the notation in [12] we decompose $D_{h}=D^{\prime}_{h}+d^{\prime\prime}$ into its $(1,0)$ and $(0,1)$ parts. The curvature of this connection is given by $R_{h}=D_{h}\wedge D_{h}$ and is an ${\rm End}E$-valued form of type $(1,1)$, usually called the Chern or the hermitian curvature of $h$. Hence, $R_{h}$ is an element in $A^{1,1}({\rm End}E)$. Using the above decomposition in types of $D_{h}$ we have
@@
Given an $h$ in $E$, the pair $(E,h)$ is called a holomorphic hermitian vector bundle and it is common to say that $D_{h}$ and $R_{h}$ are the Chern connection and curvature of $(E,h)$, respectively. Clearly, we can consider $C^{\infty}$ complex subbundles of $E$ and not only holomorphic subbundles. In particular, if $E^{\prime}\subset E$ is a $C^{\infty}$ complex subbundle we define the orthogonal complement $E^{\prime\prime}$ of $E^{\prime}$ with respect to $h$, which is also a $C^{\infty}$ subbundle of $E$. The bundles $E^{\prime}$ and $E^{\prime\prime}$ are not necessarily holomorphic and hence they give a $C^{\infty}$ orthogonal decomposition $E=E^{\prime}\oplus E^{\prime\prime}$. However, there exists a proposition in complex geometry that guarantees that the above decomposition is also holomorphic if certain invariant condition holds. To be precise one has the following result.
@@
([12], Prop. (4.18), p. 13). Let $(E,h)$ be a holomorphic hermitian vector bundle and $D_{h}$ the Chern connection of it. Let $E^{\prime}\subset E$ be a $D_{h}$-invariant $C^{\infty}$ complex subbundle and $E^{\prime\prime}$ the orthogonal complement of $E^{\prime}$ with respect to $h$. Then both $E^{\prime}$ and $E^{\prime\prime}$ are $D_{h}$-invariant holomorphic subbundles of $E$ and they give a holomorphic orthogonal decomposition:
@@
As it is well known, if $S\subset E$ is a rank $p$ holomorphic subbundle we have the following short exact sequence
@@
where $Q=E/S$ is a rank $r-p$ holomorphic vector bundle. Let $h$ be an hermitian metric in $E$ and $S^{\bot}$ the orthogonal complement of $S$ with respect to $h$, then $S^{\bot}$ is a $C^{\infty}$ complex vector bundle (not necessarily a holomorphic one) and hence $E=S\oplus S^{\bot}$ is –a priori– a $C^{\infty}$ complex orthogonal decomposition. As $C^{\infty}$ complex vector bundles $Q\cong S^{\bot}$ and the restriction of $h$ to $S^{\bot}$ induces –via this isomorphism– an hermitian metric $h_{Q}$ in $Q$. Therefore, if $h_{S}$ denotes the restriction of $h$ to $S$, the pairs $(S,h_{S})$ and $(Q,h_{Q})$ are both holomorphic hermitian vector bundles.
@@
([12], Prop. (6.4), p. 20). Let $(E,h)$ be a holomorphic hermitian vector bundle, where $D_{h}$ is the Chern connection of it and let $S\subset E$ be a holomorphic subbundle. Then, in the decomposition (1.3), the operator $D_{h,S}$ is the Chern connection of $(S,h_{S})$ and $A_{h}$ is a $(1,0)$-form with values in ${\rm Hom}(S,S^{\bot})$, i.e., $D_{h,S}=D_{h_{S}}$ and $A_{h}\in A^{1,0}({\rm Hom}(S,S^{\bot}))$.
@@
([12], Prop. (6.6), p. 21). Let $(E,h)$, $D_{h}$ and $S$ as in Proposition 2 and $Q$ the holomorphic bundle given by (1.2). Then the operator $D_{h,Q}$ defined in (1.5) is the Chern connection of $(Q,h_{Q})$ and $B_{h}$ is a $(0,1)$-form with values in ${\rm Hom}(S^{\bot},S)$, i.e., $D_{h,Q}=D_{h_{Q}}$ and $B_{h}\in A^{0,1}({\rm Hom}(S^{\bot},S))$. Moreover, for any $\xi\in A^{0}(S)$ and $\eta\in A^{0}(S^{\bot})$ we have
@@
The form $A_{h}$ in Proposition 2 is called the second fundamental form of $S$ in $(E,h)$ and again –via the isomorphism– one has $A_{h}\in A^{1,0}({\rm Hom}(S,Q))$ and $B_{h}\in A^{0,1}({\rm Hom}(Q,S))$. The expression (1.6) says that $A_{h}$ is the adjoint of $-B_{h}$ with respect to $h$. Now, as a consequence of the Leibniz rule, (1.3) and (1.4) can be naturally extended to decompositions for $s$-forms evaluated on $S$ and $S^{\bot}$, respectively. [^1] Notice that if $\lambda\in A^{s}$, then $D_{h}(\lambda\xi)=(d\lambda)\xi+D_{h}\xi\wedge\lambda=D_{h,S}(\lambda\xi)+A_{h}\wedge(\lambda\xi)\,$. Therefore, for any $\xi\in A^{0}(S)$ we have
@@
([12], Prop. (6.14), p. 23). Let $(E,h)$, $D_{h}$ and $S$ as in Proposition 2. If the second fundamental form $A_{h}$ of $S$ in $(E,h)$ vanishes identically, then the orthogonal complement $S^{\bot}\subset E$ is a holomorphic subbundle and
@@
We would like to emphasize that propositions 2 and 3 have been already extended to Higgs bundles and appear –although using a local component notation– in the form of Proposition 2.3 in [2]. However, as far as the authors know, extensions of propositions 1 and 4 have not been explored yet in literature. As we already mentioned, it is the main purpose of this article to study in detail extensions of all these propositions to Higgs bundles.

### v2

@@ Abstract.
A hermitian Higgs bundle is a triple $({\mathfrak{E}},h)=(E,\Phi,h)$, where ${\mathfrak{E}}=(E,\Phi)$ is a Higgs bundle and $(E,h)$ is a holomorphic hermitian vector bundle. It is well-known that several results on holomorphic vector bundles extend to the Higgs bundles setting, although this is not always the case. In this article we show that some classical propositions, involving orthogonal decompositions of holomorphic hermitian vector bundles and the second fundamental form of its holomorphic subbundles, can be extended to hermitian Higgs bundles. The extended propositions concerning orthogonal decompositions have immediate applications in Higgs bundles, and we mention some of these throughout the article. Moreover, the extended propositions concerning the second fundamental form are generalizations of previously known results on Higgs bundles. In particular, here we include alternative proofs of these extended propositions without using local computations. Finally, as an application of the above results and due to the lack of a certain parallelism condition, we show that a classical theorem concerning the Kobayashi functional for holomorphic vector bundles does not admit a straightforward extension to Higgs bundles.
@@ 1. Introduction
Roughly speaking a Higgs bundle is a holomorphic vector bundle, together with a section of a certain associated holomorphic vector bundle. Since its introduction by Hitchin [10] and Simpson [17, 18], these geometric objects have played an important role in complex geometry as well as in mathematical physics. For instance, from the viewpoint of geometry, several results on holomorphic bundles concerning special metrics and connections can be extended to Higgs bundles [2, 3, 4, 5, 6, 14, 16]. Similarly, from the viewpoint of physics, Higgs bundles are closely related to dimensional reduction. In fact, in the pioneering work [10], Higgs bundles arise as geometric objects associated to Hitchin’s equations (a two dimensional reduction of the self-dual Yang-Mills equations in four dimensions). On the other hand, Hitchin’s equations can also be seen as the lowest dimensional case of another set of equations introduced by Ward [21] and which are a dimensional reduction of Yang-Mills equations in higher dimensions. The equations introduced by Ward are known as $2k$-Hitchin’s equations and can be directly related to Higgs bundles [7]. It is important to mention that from the very beginning, Higgs bundles are associated to another set of equations, usually known as the Hermitian-Yang-Mills equations [17]. Higgs bundles are to the Hermitian-Yang-Mills equations as holomorphic vector bundles are to the Hermitian-Einstein equations [13, 19]. Today, Higgs bundles and all the above equations are of major interest in geometry and physics and there exists a very ample literature related to the topic, see for instance [1, 12, 20, 22]. In particular, Higgs bundles are geometric objects of relevance in the foundational work of Kapustin and Witten [11].
@@
This article is organized as follows. In Section 2 we review part of the elementary theory on holomorphic bundles that will be important later on. In particular, in this section we explicitly recall and compile some classical results on holomorphic hermitian vector bundles contained within [13], which are otherwise distributed throughout different chapters. In this section we also establish part of the notation and terminology that will be used throughout the article. In Section 3 we summarize some basic concepts and properties on Higgs bundles. We also mention two lemmas concerning short exact sequences and invariance conditions on Higgs bundles that will be relevant in the remaining part of the article. In this section we also revisit some propositions on Higgs bundles [2], which are extensions of well-known results associated to the second fundamental form of holomorphic subbundles of holomorphic hermitian vector bundles [13]. We include here alternative proofs of these results that do not use local computations. In Section 4 we establish a couple of propositions concerning orthogonal decompositions of hermitian Higgs bundles. In particular, we include proofs of these results that rely heavily on classical results on holomorphic hermitian vector bundles as well as standard definitions on Higgs bundles. We then show immediate applications of these established results. More specifically, we show that by utilizing the aforementioned lemmas and by introducing the notion of $C^{\infty}$ Higgs orthogonal complement, the extended results on Higgs bundles involving orthogonal decompositions imply two corollaries, which can be established by referring exclusively to objects in the Higgs category. Finally, in Section 5 we briefly review some properties of the Kobayashi functional which are included in [6] and we show that, due to the lack of a parallelism condition of the Hitchin-Simpson connection, a classical theorem concerning the Kobayashi functional for holomorphic vector bundles does not admit a straightforward extension to Higgs bundles.
@@
The main purpose of this article is to establish a couple of propositions as well as to revisit two other results on hermitian Higgs bundles [2]. The first couple of propositions are natural extensions of well-known results on orthogonal decompositions of holomorphic hermitian vector bundles. These classical results are included in the form of propositions 4.18 and 6.14 in the first chapter of the celebrated textbook of Kobayashi [13]. There are two other classical results, closely related to the previous propositions and which appear in the same reference as propositions 6.4 and 6.6. Extensions of these latter results to hermitian Higgs bundles have been previously studied in [2]. However, as we will see, they are are only special cases of more general propositions. For the reader’s convenience, in the following section we review some elementary definitions, as well as summarize the classical results on holomorphic vector bundles that are strictly necessary for understanding the article (for more details the reader can see [8, 9, 13]).
2. Preliminaries
@@
Throughout this article, $M$ is an $n$-dimensional compact Kähler manifold with Kähler form $\omega$ and $E$ is a rank $r$ holomorphic vector bundle over it. We denote by $\Omega^{1,0}$ and $\Omega^{0,1}$ the holomorphic cotangent bundle of $M$ and its complex conjugate bundle and by $\Omega^{p,q}$, with $0\leq p,q\leq n$, the $C^{\infty}$ complex vector bundle on $M$ obtained by taking wedge products ($p$ and $q$ times) of $\Omega^{1,0}$ and $\Omega^{0,1}$. If $1\leq s\leq 2n$ we have $\Omega^{s}=\oplus_{p+q=s}\Omega^{p,q}$ and we denote by $A^{s}$ and $A^{p,q}$ (resp. $A^{s}(E)$ and $A^{p,q}(E)$) the spaces of $C^{\infty}$ complex $s$-forms and $(p,q)$-forms on $M$ (resp. forms on $M$ with coefficients in $E$), i.e., they are the spaces of $C^{\infty}$ sections of $\Omega^{s}$ and $\Omega^{p,q}$ (resp. of $\Omega^{s}\otimes E$ and $\Omega^{p,q}\otimes E$). As it is known, any hermitian metric $h$ in $E$ determines a unique connection $D_{h}$ in $E$, usually known as the Chern or the hermitian connection of $h$. Following the notation in [13] we decompose $D_{h}=D^{\prime}_{h}+d^{\prime\prime}$ into its $(1,0)$ and $(0,1)$ parts. The curvature of this connection is given by $R_{h}=D_{h}\wedge D_{h}$ and is an ${\rm End}E$-valued form of type $(1,1)$, usually called the Chern or the hermitian curvature of $h$. Hence, $R_{h}$ is an element in $A^{1,1}({\rm End}E)$. Using the above decomposition in types of $D_{h}$ we have
@@
Associated to the curvature $R_{h}$, one has the Chern mean curvature $K_{h}$, which is the element in $A^{0}({\rm End}E)$ defined by [^1] Some sources [2, 16] equivalently define the mean curvature operator as $K_{h}=i\Lambda_{\omega}R_{h}$, where $\Lambda_{\omega}$ is the dual of the Lefschetz operator $L=\omega\wedge\bullet$ (see [13, p. 99] for an explanation of this equivalence).
(2.1) $inR_{h}\wedge\omega^{n-1}=K_{h}\,\omega^{n}.$
A hermitian metric $h$ is said to be Hermitian-Yang-Mills if $K_{h}=cI$, where $I$ denotes the identity endomorphism of $E$ and $c$ is a constant, which is indeed fixed by invariants of the bundle (see [13], p. 178). If $h$ is Hermitian-Yang-Mills, the pair $(E,h)$ is called an Hermitian-Yang-Mills vector bundle. [^2] Sometimes the terminology Hermitian-Einstein metric is also used, the latest one was indeed introduced by Kobayashi [13] as a generalization to holomorphic vector bundles of the notion of Kähler-Einstein metric for complex Kähler manifolds.
Given an $h$ in $E$, the pair $(E,h)$ is called a holomorphic hermitian vector bundle and it is common to say that $D_{h}$ and $R_{h}$ are the Chern connection and curvature of $(E,h)$, respectively. Clearly, we can consider $C^{\infty}$ complex subbundles of $E$ and not only holomorphic subbundles. In particular, if $E^{\prime}\subset E$ is a $C^{\infty}$ complex subbundle we define the orthogonal complement $E^{\prime\prime}$ of $E^{\prime}$ with respect to $h$, which is also a $C^{\infty}$ subbundle of $E$. The bundles $E^{\prime}$ and $E^{\prime\prime}$ are not necessarily holomorphic and hence they give a $C^{\infty}$ orthogonal decomposition $E=E^{\prime}\oplus E^{\prime\prime}$. However, there exists a proposition in complex geometry that guarantees that the above decomposition is also holomorphic if a certain invariant condition holds. To be precise one has the following result.
@@
([13], Prop. 4.18, p. 13). Let $(E,h)$ be a holomorphic hermitian vector bundle and $D_{h}$ its Chern connection. Let $E^{\prime}\subset E$ be a $D_{h}$-invariant $C^{\infty}$ complex subbundle and $E^{\prime\prime}$ the orthogonal complement of $E^{\prime}$ with respect to $h$. Then both $E^{\prime}$ and $E^{\prime\prime}$ are $D_{h}$-invariant holomorphic subbundles of $E$ and they give a holomorphic orthogonal decomposition:
@@
There exists an important application of Proposition 1 involving the holonomy of the Chern connection of $(E,h)$ and which we briefly summarize here (for more details see [13], p. 107). Let $\Psi(x)$ be the holonomy group of $D_{h}$ with reference point $x\in M$ and let
(2.3) $E_{x}=E_{x}^{0}\oplus E_{x}^{1}\oplus\cdots\oplus E_{x}^{k}$
be the orthogonal decomposition of the fibre $E_{x}$ such that $\Psi(x)$ acts trivially on $E_{x}^{0}$ and irreducibly on each $E_{x}^{1},...,E_{x}^{k}$. Of course $E_{x}^{0}$ may be the trivial vector space. Since $D_{h}$ is the Chern connection, $h$ is parallel with respect to it. Therefore, by taking a parallel displacement of (2.3) we obtain an orthogonal decomposition
(2.4) $E=E^{0}\oplus E^{1}\oplus\cdots\oplus E^{k}$
where $E^{0}$ is trivial as a hermitian vector bundle. Since $E^{0},...,E^{k}$ are all $D_{h}$-invariant $C^{\infty}$ complex subbundles of $E$, Proposition 1 implies that (2.4) is also holomorphic, i.e., it is a holomorphic orthogonal decomposition of $E$.
As it is well-known, if $S\subset E$ is a rank $p$ holomorphic subbundle we have the following short exact sequence
@@
where $Q=E/S$ is a rank rpr-p holomorphic vector bundle. Let $h$ be a hermitian metric in $E$ and $S^{\bot}$ the orthogonal complement of $S$ with respect to $h$, then $S^{\bot}$ is a $C^{\infty}$ complex vector bundle (not necessarily a holomorphic one) and hence $E=S\oplus S^{\bot}$ is –a priori– a $C^{\infty}$ complex orthogonal decomposition. As $C^{\infty}$ complex vector bundles $Q\cong S^{\bot}$ and the restriction of $h$ to $S^{\bot}$ induces –via this isomorphism– a hermitian metric $h_{Q}$ in $Q$. Therefore, if $h_{S}$ denotes the restriction of $h$ to $S$, the pairs $(S,h_{S})$ and $(Q,h_{Q})$ are both holomorphic hermitian vector bundles.
@@
([13], Prop. 6.4, p. 20). Let $(E,h)$ be a holomorphic hermitian vector bundle and $D_{h}$ its Chern connection. Let $S\subset E$ be a holomorphic subbundle. Then, in the decomposition (2.6), the operator $D_{h,S}$ is the Chern connection of $(S,h_{S})$ and $A_{h}$ is a $(1,0)$-form with values in ${\rm Hom}(S,S^{\bot})$, i.e., $D_{h,S}=D_{h_{S}}$ and $A_{h}\in A^{1,0}({\rm Hom}(S,S^{\bot}))$.
@@
([13], Prop. 6.6, p. 21). Let $(E,h)$, $D_{h}$ and $S$ as in Proposition 2 and $Q$ the holomorphic bundle given by (2.5). Then the operator $D_{h,Q}$ defined in (2.8) is the Chern connection of $(Q,h_{Q})$ and $B_{h}$ is a $(0,1)$-form with values in ${\rm Hom}(S^{\bot},S)$, i.e., $D_{h,Q}=D_{h_{Q}}$ and $B_{h}\in A^{0,1}({\rm Hom}(S^{\bot},S))$. Moreover, for any $\xi\in A^{0}(S)$ and $\eta\in A^{0}(S^{\bot})$ we have
@@
The form $A_{h}$ in Proposition 2 is called the second fundamental form of $S$ in $(E,h)$ and again –via the isomorphism– one has $A_{h}\in A^{1,0}({\rm Hom}(S,Q))$ and $B_{h}\in A^{0,1}({\rm Hom}(Q,S))$. The expression (2.9) says that $A_{h}$ is the adjoint of $B_{h}$ with respect to $h$. Now, as a consequence of the Leibniz rule, (2.6) and (2.7) can be naturally extended to decompositions for $s$-forms evaluated on $S$ and $S^{\bot}$, respectively. [^3] Notice that if $\lambda\in A^{s}$, then $D_{h}(\lambda\xi)=(d\lambda)\xi+D_{h}\xi\wedge\lambda=D_{h,S}(\lambda\xi)+A_{h}\wedge(\lambda\xi)\,$. Therefore, for any $\xi\in A^{0}(S)$ we have
@@
([13], Prop. 6.14, p. 23). Let $(E,h)$, $D_{h}$ and $S$ as in Proposition 2. If the second fundamental form $A_{h}$ of $S$ in $(E,h)$ vanishes identically, then the orthogonal complement $S^{\bot}\subset E$ is a holomorphic subbundle and
@@
On the other hand, let us denote by ${\rm Herm}^{+}E$ the space of hermitian metrics in the holomorphic vector bundle $E$. If $h\in{\rm Herm}^{+}E$ is considered as a variable and $K_{h}$ is the curvature defined in (2.1), Kobayashi introduces the functional $J:{\rm Herm}^{+}E\longrightarrow{\mathbb{R}}$ defined by
(2.12) $J(h)=\frac{1}{2}\int_{M}\lvert K_{h}\lvert^{2}\omega^{n}=\frac{n!}{2}\lVert K_{h}\lVert^{2}\,$
where $\lvert\cdot\lvert$ and $\lVert\cdot\lVert$ denote the usual pointwise and $L^{2}$-norms in $A^{0}({\rm End}E)$, respectively (see [13] or [6, 7] for details). Hence, up to a multiplicative constant, $J$ is the energy functional for the Chern mean curvature. As it is well-known, the functional (2.12) can be considered as the Yang-Mills functional [^4] Strictly speaking, the Yang-Mills functional $I$ is the energy functional of the Chern curvature. However, the difference I(h)J(h)I(h)-J(h) does not depend on $h$ and is a topological term involving the first and second Chern classes of $E$. See [13], p. 111 for details. and there are some classical results associated to $J$. In particular, a hermitian metric $h$ is a minimum of $J$ if and only if it is Hermitian-Yang-Mills. More in general one has the following result.
Theorem 1.
([13], Thm. 3.21, p. 110). Let $E$ be a holomorphic vector bundle. A hermitian metric $h$ is a critical point of the functional (2.12) if and only if the Chern mean curvature is parallel with respect to the Chern connection defined by $h$, i.e., if and only if
(2.13) $D_{h}K_{h}=0\,.$
The parallelism condition (2.13), together with the parallelism of $h$ with respect to the Chern connection, has very important consequences for the holomorphic bundle $E$ and functional $J$ (see [13], p. 111 for details). In fact, assume that $h$ is a critical point of $J$. Then, using (2.13) we can diagonalize $K_{h}$ and obtain a decomposition of $E$ in terms of the $C^{\infty}$ subbundles. Taking a further decomposition of it using the holonomy group of $D_{h}$, we obtain a holomorphic orthogonal decomposition as in (2.4), where this time $E^{0}$, $E^{1},...,E^{k}$ are holomorphic eigenbundles of $K_{h}$ with constants $0,c_{1},...,c_{k}$. In summary, from the above one has the following result.
Theorem 2.
([13], Thm. 3.27, p. 111). Let $E$ be a holomorphic vector bundle. Assume that $h$ is a critical point of the functional (2.12) and let
(2.14) $E=E^{0}\oplus E^{1}\oplus\cdots\oplus E^{k}\,$
be the holomorphic orthogonal decomposition of $(E,h)$ given by (2.4). Let $h_{0},h_{1},...,h_{k}$ be the restrictions of $h$ to $E^{0},E^{1},...,E^{k}$, respectively. Then, the pairs $(E^{0},h_{0}),(E^{1},h_{1}),...,(E^{k},h_{k})$ are all Hermitian-Yang-Mills vector bundles with constants $0,c_{1},...,c_{k}$.
We would like to emphasize that propositions 2 and 3 have been already extended to Higgs bundles and appear –although using a local component notation– in the form of Proposition 2.3 in [2]. However, as we will explain in the following section, these results are just special cases of two more general propositions on Higgs bundles. Also, Theorem 1 has been extended to Higgs bundles in [6]. However, as far as the authors know, extensions of propositions 1 and 4 and Theorem 2 have not been explored yet in the literature. As previously mentioned, the main purpose of this article is to study in detail possible extensions of these results to Higgs bundles. We will address all these issues in the last two sections of the article.